# Inverse 3D microscopy rendering for cell shape inference with active mesh

### v2

@@ 3.2 Point spread function model
$h(\mathbf{z})=\dfrac{\mathrm{e}^{-\frac{1}{2}\mathbf{z}^{T}\Sigma^{-1}\mathbf{z}}}{\sqrt{(2\pi)^{n}\det\Sigma}}\;,\;\mathbf{z}\in\mathbb{R}^{3}.$ (2)
The FT $\hat{h}$ of $h$ is given by: $\hat{h}(\boldsymbol{\xi})=\mathrm{e}^{-\frac{1}{2}\boldsymbol{\xi}^{T}\Sigma^{-1}\boldsymbol{\xi}},$ where $\boldsymbol{\xi}\in\mathbb{R}^{3}$ is the wavevector. In this framework, inferring the PSF consists of optimizing the coefficients of the covariance matrix $\Sigma$, which is symmetric positive semi-definite and directly controls the level of blur in the rendered image. More photorealistic PSFs based on differentiable physical models could also be incorporated [23, 24], allowing the learned parameters to correspond to real optical properties, such as refractive indices or optical aberrations expressed using Zernike polynomials [40].
@@ 3.3 Geometric models of biological objects
Extracting geometry from volumetric images involves approximating biological objects using discrete $ND$ representations ($N=0,1,2,3$) embedded in 3D space. These objects vary widely in size and topology. Here, we focus exclusively on cell membranes, approximating them as 2D surfaces efficiently modeled with triangle meshes, but note that extensions to point-like, filaments or bulk structures would also be possible.This choice incorporates prior knowledge of structural topology: as shown in Section 6.4, early embryos consist of cells forming bounded regions, which can be represented as a single non-manifold multimaterial mesh [13, 49], while capturing cell-cell interfaces by doubling intensity due to overlapping adhesive membranes.
@@ 4 Fourier transform of triangulated surfaces
In this section, we derive an explicit expression for the Fourier Transform (FT) of arbitrary 2D surfaces embedded in $\mathbb{R}^{3}$ and provide closed-form formulas for both the FT and its gradient with respect to vertex positions in the specific case of a triangulated mesh.
Figure 3: Pipeline Overview. The microscopy renderer generates a 3D artificial fluorescence image from a triangle mesh representing a biological object and a parameterized PSF simulating the microscope’s optics. Gradients with respect to the mesh and PSF parameters are computed via backpropagation, enabling optimization to align the output with real microscopy data.
@@ 4.1 Surfaces as spatial Dirac distributions
The density of fluorophores $u_{\Lambda}$ is equal to 1 on surfaces $\Lambda$ (2 in case of double membranes), and 0 everywhere else.
Considering a surface parametrized by a function $\boldsymbol{\Lambda}:[0,1]^{2}\to[0,1]^{3}$, its spatial density is given by:
$u_{\Lambda}=\dfrac{1}{|\Lambda|}\int_{0}^{1}\int_{0}^{1}\delta_{\Lambda}(x,y)a_{\Lambda}(x,y)\mathrm{d}x\mathrm{d}y,$ (3)
where $\delta_{\Lambda}(x,y)$ is the Dirac function, $a_{\Lambda}$ is the surface element given by $a_{\Lambda}(x,y)=\|\partial_{x}\Lambda\times\partial_{y}\Lambda\|$, and $|\Lambda|$ is the total area of the surface $\Lambda$. We normalize by the total surface area to ensure a total density $1$. When dealing with several surfaces with different densities, one can simply used a weighted sum or integral of $u_{\Lambda}$.
For all $\mathbf{z}\in\mathbb{R}^{3}$, the FT of the Dirac distribution at z is given by $\hat{\delta}_{\mathbf{z}}(\boldsymbol{\xi})=e^{-i\mathbf{z}\cdot\boldsymbol{\xi}},\,\boldsymbol{\xi}\in\mathbb{R}^{3}.$ By linearity, the FT of $u_{\Lambda}$ reads:
$\hat{u}_{\Lambda}(\boldsymbol{\xi})=\dfrac{1}{|\Lambda|}\int_{0}^{1}\int_{0}^{1}a_{\Lambda}(x,y)e^{-i\boldsymbol{\Lambda}(x,y)\cdot\boldsymbol{\xi}}\mathrm{d}x\mathrm{d}y.$ (4)
In practice, we discretize here $\Lambda$ using triangle meshes. By linearity of the FT, this amounts to compute the FT of each triangle of the surface and sum their contributions.
4.1.1 Density for a triangle mesh
@@
$\hat{u}_{\Lambda}(\boldsymbol{\xi})=\frac{\underset{\mathcal{T}\in\Lambda}{\sum}\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})}{|\Lambda|}=\frac{\underset{\mathcal{T}\in\Lambda}{\sum}\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})}{\underset{\mathcal{T}\in\Lambda}{\sum}A_{\mathcal{T}}}.$ (5)
The gradient of the FT with respect to a vertex v reads:
@@ 4.1.2 Fourier transform and gradients of a triangle
We consider a triangle $\mathcal{T}$ defined by its vertices $(\textbf{v}_{1},\textbf{v}_{2},\textbf{v}_{3})$ and compute $u_{\Lambda}$ and its spatial derivatives. For convenience, we set $\textbf{v}_{4}\!=\!\textbf{v}_{1}$ and similarly by permutation $\textbf{v}_{0}\!=\!\textbf{v}_{3}$. For $p\!=\!1\ldots 3$, we define $p^{-}\!=\!p\!-\!1$ and $p^{+}\!=\!p\!+\!1$ and denote by $\textbf{e}_{p}=\textbf{v}_{p^{-}}\!-\!\textbf{v}_{p^{+}}$ the opposite edge to $\textbf{v}_{p}$ and by $l_{p}=|\textbf{e}_{p}|$ its length. $A_{\mathcal{T}}\!=\!\dfrac{|\textbf{e}_{3}\times\textbf{e}_{1}|}{2}$ denotes the area of the triangle, and $\textbf{N}_{\mathcal{T}}\!=\!\dfrac{\textbf{e}_{3}\times\textbf{e}_{1}}{2A_{\mathcal{T}}}$ denotes its unit normal. For $p\!=1\!\ldots 3$, we define $\textbf{w}_{p}=\textbf{e}_{p}\times\textbf{N}_{\mathcal{T}}$, the non-normalized outward normal to $\mathcal{T}$ on the edge $\textbf{e}_{p}$. This allows us expressing the gradient of the triangle area with respect to each vertex position as $\dfrac{\partial A_{\mathcal{T}}}{\partial\textbf{v}_{p}}=-\dfrac{\textbf{w}_{p}}{2}$, for $p=1\ldots 3$.
Expressing the FT of the density on a triangle as $\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})=\int_{\mathcal{T}}\mathrm{e}^{-i\mathbf{z}\cdot\boldsymbol{\xi}}\mathrm{d}s(\mathbf{z})$, then for all $\boldsymbol{\xi}\in\mathbb{R}^{3}$, we have:
@@
$\text{with}\quad f_{\mathcal{T}}(\boldsymbol{\xi})=\sum_{p=1}^{3}\frac{\mathrm{e}^{-i\mathbf{z}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})}.$ (8)
@@
$\begin{split}\text{with}\,\dfrac{\partial f_{\mathcal{T}}(\boldsymbol{\xi})}{\partial\textbf{v}_{p}}=&\boldsymbol{\xi}\left[\dfrac{\text{e}^{-\text{i}\mathbf{z}_{p^{+}}\cdot\boldsymbol{\xi}}}{(\textbf{v}_{p^{-}}\cdot\boldsymbol{\xi})^{2}(\textbf{v}_{p}\cdot\boldsymbol{\xi})}-\dfrac{\text{e}^{-\text{i}\mathbf{z}_{p^{-}}\cdot\boldsymbol{\xi}}}{(\textbf{v}_{p}\cdot\boldsymbol{\xi})(\textbf{v}_{p^{+}}\cdot\boldsymbol{\xi})^{2}}\right.\\ &\left.-\dfrac{\text{i}\text{e}^{-\text{i}\mathbf{z}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{v}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{v}_{p^{+}}\cdot\boldsymbol{\xi})}+\dfrac{\text{e}^{-\text{i}\mathbf{z}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{v}_{p^{-}}\cdot\boldsymbol{\xi})^{2}(\textbf{v}_{p^{+}}\cdot\boldsymbol{\xi})}\right.\\ &\left.-\dfrac{\text{e}^{-\text{i}\mathbf{z}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{v}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{v}_{p^{+}}\cdot\boldsymbol{\xi})^{2}}\right].\end{split}$ (10)
@@ 4.1.3 Numerical approximations to prevent divergence
$\begin{array}[]{ll}g(s,t,u)\!=\!\dfrac{-\,\mathrm{e}^{is}}{(s\!-\!t)(s\!-\!u)}\!-\!\dfrac{\mathrm{e}^{it}}{(t\!-\!u)(t\!-\!s)}\!-\!\dfrac{\mathrm{e}^{iu}}{(u\!-\!s)(u\!-\!t)}.\end{array}$ (11)
When two values $(a,b)$ among $(s,t,u)$ satisfy $|a\!-\!b|<\epsilon$, the denominator in Eq. (11) approaches zero, leading to divergence. To prevent this, we derive exact expressions for $g(s,t,u)$ in the special cases where $t=u,t=s,u=s$, or $u=s=t$ These alternative expressions replace the original formulation in Eq. (11) whenever $|a\!-\!b|<\epsilon$, ensuring numerical stability:
@@
$=i\dfrac{\mathrm{e}^{-it}}{t-u}+\dfrac{\mathrm{e}^{-it}}{(t-u)^{2}}-\dfrac{\mathrm{e}^{-iu}}{(t-u)^{2}}.$ (12a)
$g(u,u,u)$ $=\frac{\mathrm{e}^{-iu}}{2}.$ (12b)
@@ 5.1 Acceleration
Computing the artificial image requires evaluating the FT $\hat{u}_{\Lambda}$ for each of the $N$ voxels times the number $n_{t}$ of triangles in the mesh, resulting in a runtime complexity of $\mathcal{O}(N\cdot n_{t})$. The gradient of the mesh FT involves an $\mathcal{O}(1)$ sum per voxel for each of the $n_{v}$ vertices, leading to a total complexity of $\mathcal{O}(N\cdot n_{v})$. In practice, a confocal microscopy image of size $500^{3}$ contains 125 million voxels, while a reasonable mesh typically consists of $10^{3-4}$ vertices and triangles. Without proper parallelization, these computations become prohibitive. The high computational cost has been a key limitation preventing the widespread use of spectral methods in 3D, despite their promising applications. [30].
Figure 5: Benchmarking of Mesh FT Computation. (Left) For a small 80-triangle sphere, our CUDA implementation significantly improves speed and memory efficiency. (Right) In a $[100]^{3}$, the forward pass scales linearly with the number of triangles.
@@
GPU parallelization The grid-based structure of a 3D image naturally lends itself to massively parallel computations on graphical processing units (GPUs). To leverage this, we provide a custom CUDA implementation for both the forward and backward passes. As shown in Figure 5, our implementation achieves a $\approx 10^{3}$ speedup on a single NVIDIA V100 GPU compared to a vectorized CPU implementation, while also enabling computations on significantly larger box sizes before encountering overflow.
Narrow-band approximation in the frequency domain As previously stated, the Fourier Transform (FT) of the image is given by the element-wise product of the FT of the mesh and the PSF: $\hat{u}_{\alpha}=\hat{u}_{\Lambda}\cdot\hat{h}$. In practice, the PSF acts as a low-pass filter, emphasizing low spatial frequencies of the mesh. A blurred image corresponds to a sparse PSF, with its highest amplitudes concentrated near zero frequency, thereby suppressing fine mesh details. The more resolved the original biological image, the more frequency components of the PSF are required for accurate rendering. For any given spatial frequency $\boldsymbol{\xi}$, if $\hat{h}(\boldsymbol{\xi})\approx 0$, then computing $\hat{u}_{\Lambda}(\boldsymbol{\xi})$ is unnecessary, as it contributes negligibly to $\hat{u}_{\alpha}(\boldsymbol{\xi})$. To exploit this, we apply a frequency-domain cutoff, computing the FT of the mesh only for frequencies where the PSF exceeds $1\%$ of its maximum value. For sparse PSFs, this spectral narrow-band method reduces computational cost by several orders of magnitude.
@@
A staggered optimization scheme Despite significant speed improvements, computing the mesh FT and its gradient remains the main computational bottleneck of our pipeline. We offer two approaches for computing the FT: a fast, approximate method using the narrow-band approach and a slower, exact method performing the full computation. The forward pass is approximately 10 times slower than the backward pass, making it crucial to apply the narrow-band method in the backward computation whenever possible. To address this, we decompose the coupled optimization problem into two staggered sub-problems, iterating them sequentially. (a) Shape optimization: To optimize the shape, an approximate computation of the mesh FT suffices. We use the narrow-band method to efficiently compute the FT and perform an optimization step for the vertex positions while keeping the PSF fixed.. (b) PSF optimization: Optimizing the PSF requires computing the mesh FT across all available frequencies, avoiding bias from the narrow-band threshold, which depends on the current PSF values. Thus, we compute the exact FT of the mesh and perform an optimization step for the PSF parameters while keeping the shape fixed.
5.2 Shape optimization step
We optimize shapes by minimizing a modified weighted L2 norm $\Phi=\left<(u-m)^{2},m\right>$ that removes external influence from dark regions of the original image $m$. Inspired by other works in differentiable rendering [58], we optimize vertex positions using AdamUniform and regularize gradient values by adding a diffusion term: $\textbf{v}\leftarrow x-\eta(\textbf{I}+\lambda\textbf{L})^{-2}\dfrac{\partial\Phi}{\partial\textbf{v}}$, where $\textbf{L}\in\mathbb{R}^{n_{v}\times n_{v}}$ is a graph Laplacian defined on our triangle mesh, and $\lambda=50$. Computing explicitly the inverse $(\textbf{I}+\lambda\textbf{L})^{-2}$ requires to use dense algebra, which can lead to prohibitive memory usage when the number of vertices $n_{v}$ of the mesh grows. Instead, we solve the equivalent problem of finding the solution A to $(\textbf{I}+\lambda\textbf{L})^{-2}\text{A}=\dfrac{\partial\Phi}{\partial\textbf{v}}$ with a sparse Cholesky decomposition [56, 58].
Figure 6: Shape inference from 3D artificial images. Starting from spherical or toroidal meshes, the optimization process reconstructs meshes that closely align with the original shapes (see also supplementary videos 1–4).

### v3

@@ 3.2 Point spread function model
$h(\mathbf{x})=\dfrac{\mathrm{e}^{-\frac{1}{2}\mathbf{x}^{T}\Sigma^{-1}\mathbf{x}}}{\sqrt{(2\pi)^{3}\det\Sigma}}\;,\;\mathbf{x}\in\mathbb{R}^{3}.$ (2)
The FT $\hat{h}$ of $h$ is given by: $\hat{h}(\boldsymbol{\xi})=\mathrm{e}^{-\frac{1}{2}\boldsymbol{\xi}^{T}\Sigma^{-1}\boldsymbol{\xi}},$ where $\boldsymbol{\xi}\in\mathbb{R}^{3}$ is the wavevector. In this framework, inferring the PSF consists of optimizing the coefficients of the symmetric positive semi-definite covariance matrix $\Sigma$, which directly controls the level of blur in the rendered image. More photorealistic PSFs based on differentiable physical models could also be incorporated [29, 30], allowing the learned parameters to correspond to real optical properties, such as refractive indices or optical aberrations expressed using Zernike polynomials [47].
@@ 3.3 Geometric models of biological objects
Extracting geometry from volumetric images involves approximating biological structures using discrete $N$-dimensional representations ($N=0,1,2,3$) embedded in 3D space. These structures vary in size and topology. We focus here on cell membranes, modeled as 2D surfaces using triangle meshes, though extensions to points, filaments, and volumes are possible. This choice reflects prior knowledge of structural topology: as shown in Section 6.4, early embryos comprise cells forming bounded regions, representable as a single non-manifold, multimaterial mesh [15, 56], where cell–cell interfaces are captured by doubling intensity at overlapping membranes.
@@ 4 Fourier transform of triangulated surfaces
We derive an explicit expression for the FT of arbitrary 2D surfaces embedded in $\mathbb{R}^{3}$, and provide closed-form formulas for both the FT and its gradient with respect to vertex positions in the case of a triangulated mesh.
@@ 4.1 Surfaces as spatial Dirac distributions
We model the spatial fluorophore density $u_{\Lambda}$ as a Dirac distribution supported on the surface $\Lambda$, weighted by the local surface element to reflect the actual fluorophore density along the surface.
Figure 4: Surface $\Lambda$ describing a dividing cell and its associated fluorophore density $u_{\Lambda}$ convolved with a Gaussian PSF in $\mathbb{R}^{3}$ (maximum projection).
We consider a surface parametrized by a function $\boldsymbol{\Lambda}:\,[0,1]^{2}\to[0,1]^{3},$ mapping local coordinates $(s_{1},s_{2})$, defined in a tangent orthonormal basis, to points in 3D space. The spatial fluorophore density in $\mathbb{R}^{3}$ is then defined as:
$u_{\Lambda}(\mathbf{x})=\dfrac{1}{|\Lambda|}\iint\delta_{\boldsymbol{\Lambda}(s_{1},s_{2})}(\mathbf{x})\,a_{\Lambda}(s_{1},s_{2})\,\mathrm{d}s_{1}\mathrm{d}s_{2},$ (3)
where $\delta_{\boldsymbol{\Lambda}(s_{1},s_{2})}$ denotes the Dirac delta function centered at $\boldsymbol{\Lambda}(s_{1},s_{2})$, $a_{\Lambda}\!=\!\|\partial_{s_{1}}\boldsymbol{\Lambda}\!\times\!\partial_{s_{2}}\boldsymbol{\Lambda}\|$ is the surface element. The normalization by the total surface area $|\Lambda|=\iint a_{\Lambda}(s_{1},s_{2})\mathrm{d}s_{1}\mathrm{d}s_{2}$ ensures that the total integrated image intensity is one. When dealing with several surfaces with different densities, one can use a weighted sum or integral of $u_{\Lambda}$. Cell-cell interfaces are modeled as double membranes, and their fluorophore density is therefore scaled by a factor two.
For all $\mathbf{x}\!=\!(x,y,z)\in\mathbb{R}^{3}$, the FT of the Dirac distribution at a point $\mathbf{x}$ is given by $\hat{\delta}_{\mathbf{x}}(\boldsymbol{\xi})=\mathrm{e}^{-\mathrm{i}\mathbf{x}\cdot\boldsymbol{\xi}},\,\boldsymbol{\xi}\in\mathbb{R}^{3}.$ By linearity, the FT of $u_{\Lambda}$ reads therefore:
$\hat{u}_{\Lambda}(\boldsymbol{\xi})=\dfrac{1}{|\Lambda|}\iint a_{{\Lambda}}(s_{1},s_{2})\,\mathrm{e}^{-\mathrm{i}\boldsymbol{\Lambda}(s_{1},s_{2})\cdot\boldsymbol{\xi}}\,\mathrm{d}s_{1}\mathrm{d}s_{2}.$ (4)
In practice, we discretize $\Lambda$ with triangle meshes. By linearity, this reduces to computing and summing the FT of each triangle.
4.1.1 Spatial density for a triangle mesh
@@
$\hat{u}_{\Lambda}(\boldsymbol{\xi})=\frac{\underset{\mathcal{T}\in\Lambda}{\sum}\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})}{|\Lambda|}=\frac{\underset{\mathcal{T}\in\Lambda}{\sum}\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})}{\underset{\mathcal{T}\in\Lambda}{\sum}A_{\mathcal{T}}},$ (5)
Where $A_{\mathcal{T}}$ is the area of $\mathcal{T}$. The gradient of the FT with respect to a vertex v reads:
@@ 4.1.2 Fourier transform and gradients of a triangle
We consider a triangle $\mathcal{T}$ defined by its vertices $(\textbf{v}_{1},\textbf{v}_{2},\textbf{v}_{3})$ and compute $\hat{u}_{\mathcal{T}}$ and its spatial derivatives. For convenience, we set $\textbf{v}_{4}\!=\!\textbf{v}_{1}$ and similarly by permutation $\textbf{v}_{0}\!=\!\textbf{v}_{3}$. For $p\!=\!1\ldots 3$, we define $p^{-}\!=\!p\!-\!1$ and $p^{+}\!=\!p\!+\!1$ and denote by $\textbf{e}_{p}=\textbf{v}_{p^{-}}\!-\!\textbf{v}_{p^{+}}$ the opposite edge to $\textbf{v}_{p}$ and by $l_{p}=|\textbf{e}_{p}|$ its length. $A_{\mathcal{T}}\!=\!\dfrac{|\textbf{e}_{3}\times\textbf{e}_{1}|}{2}$ denotes the area of the triangle, and $\textbf{N}_{\mathcal{T}}\!=\!\dfrac{\textbf{e}_{3}\times\textbf{e}_{1}}{2A_{\mathcal{T}}}$ denotes its unit normal. For $p\!=1\!\ldots 3$, we define $\textbf{w}_{p}=\textbf{e}_{p}\times\textbf{N}_{\mathcal{T}}$, the non-normalized outward normal to $\mathcal{T}$ on the edge $\textbf{e}_{p}$. This allows us expressing the gradient of the triangle area with respect to each vertex position as $\dfrac{\partial A_{\mathcal{T}}}{\partial\textbf{v}_{p}}=-\dfrac{\textbf{w}_{p}}{2}$, for $p=1\ldots 3$.
Expressing the FT of the density on a triangle as $\hat{u}_{\mathcal{T}}(\boldsymbol{\xi})=\int_{\mathcal{T}}\mathrm{e}^{-\mathrm{i}\mathbf{x}\cdot\boldsymbol{\xi}}\mathrm{d}s(\mathbf{x})$, then for all $\boldsymbol{\xi}\in\mathbb{R}^{3}$, we have:
@@
$\text{with}\quad f_{\mathcal{T}}(\boldsymbol{\xi})$ $=\sum_{p=1}^{3}\frac{\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})}.$ (8)
@@
$\begin{split}\text{with}\,\dfrac{\partial f_{\mathcal{T}}(\boldsymbol{\xi})}{\partial\textbf{v}_{p}}=&\boldsymbol{\xi}\left[\dfrac{\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p^{+}}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})^{2}(\textbf{e}_{p}\cdot\boldsymbol{\xi})}-\dfrac{\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p^{-}}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p}\cdot\boldsymbol{\xi})(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})^{2}}\right.\\ &\left.-\dfrac{\mathrm{i}\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})}+\dfrac{\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})^{2}(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})}\right.\\ &\left.-\dfrac{\mathrm{e}^{-\mathrm{i}\mathbf{v}_{p}\cdot\boldsymbol{\xi}}}{(\textbf{e}_{p^{-}}\cdot\boldsymbol{\xi})(\textbf{e}_{p^{+}}\cdot\boldsymbol{\xi})^{2}}\right].\end{split}$ (10)
@@ 4.1.3 Numerical approximations to prevent divergence
$\begin{array}[]{ll}g(s,t,u)\!=\!\dfrac{-\,\mathrm{e}^{is}}{(s\!-\!t)(s\!-\!u)}\!+\!\dfrac{-\,\mathrm{e}^{it}}{(t\!-\!u)(t\!-\!s)}\!+\!\dfrac{-\,\mathrm{e}^{iu}}{(u\!-\!s)(u\!-\!t)}.\end{array}$ (11)
[…]
@@
$=i\dfrac{\mathrm{e}^{-\mathrm{i}t}}{t-u}+\dfrac{\mathrm{e}^{-\mathrm{i}t}}{(t-u)^{2}}-\dfrac{\mathrm{e}^{-\mathrm{i}u}}{(t-u)^{2}}.$ (12a)
$g(u,u,u)$ $=\frac{\mathrm{e}^{-\mathrm{i}u}}{2}.$ (12b)
@@ 5.1 Acceleration
Computing the artificial image requires evaluating the FT $\hat{u}_{\Lambda}$ for each of the $N$ voxels times the number $n_{t}$ of triangles in the mesh, resulting in a runtime complexity of $\mathcal{O}(N\cdot n_{t})$. The gradient of the mesh FT involves an $\mathcal{O}(1)$ sum per voxel for each of the $n_{v}$ vertices, leading to a total complexity of $\mathcal{O}(N\cdot n_{v})$. In practice, a confocal microscopy image of size $500^{3}$ contains 125 million voxels, while a reasonable mesh typically consists of $10^{3-4}$ vertices and triangles. Without proper parallelization, these computations become prohibitive. The high computational cost has been a key limitation preventing the widespread use of spectral methods in 3D, despite their promising applications [37].
@@
GPU parallelization The grid-based structure of a 3D image naturally lends itself to massively parallel computations on graphical processing units (GPUs). To leverage this, we provide a custom CUDA implementation for both, the forward and the backward passes. As shown in Figure 6, our implementation achieves $\approx 10^{3}$ speedup on one NVIDIA V100 GPU compared to a vectorized CPU implementation, while also enabling computations on significantly larger box sizes before encountering overflow.
Figure 6: Benchmarking of Mesh FT Computation: (Left) For a 80-triangle spherical mesh, our CUDA implementation significantly improves speed and memory efficiency as function of image size. (Right) For an image of size $[100]^{3}$, the forward pass scales linearly with the number of triangles.
[…]
@@
A staggered optimization scheme Despite significant speed improvements, computing the mesh FT and its gradient remains the main computational bottleneck of our pipeline. We offer two approaches for computing the FT: a fast, approximate method using the narrow-band approach and a slower, exact method performing the full computation. The forward pass is approximately 10 times slower than the backward pass, making it crucial to apply the narrow-band method in the backward computation whenever possible. To address this, we decompose the coupled optimization problem into two staggered sub-problems, iterating them sequentially. (a) Shape optimization: To optimize the shape, an approximate computation of the mesh FT suffices. We use the narrow-band method to efficiently compute the FT and perform an optimization step for the vertex positions while keeping the PSF fixed. (b) PSF optimization: Optimizing the PSF requires computing the mesh FT across all available frequencies, avoiding bias from the narrow-band threshold, which depends on the current PSF values. Thus, we compute the exact FT of the mesh and perform an optimization step for the PSF parameters while keeping the shape fixed.
[…]